# Unifying Attention Heads and Task Vectors via Hidden State Geometry in In-Context Learning

### v1

@@ 5.2 Semantic interpretation of alignment dynamics
Figure 4: Dynamics of alignment measures and semantics of post-transition hidden states. (A) Strong correlation between output and directional alignment; (B) Surge in alignment measures concur with encoding of label-related semantics; (C) Post-transition layers, except for the last ones, filter out unrelated semantics and retain the related ones. Refer to Appendix F for more semantics decoding cases.
@@
To understand how the surge in directional alignment measures after the phase transition ultimately leads to high output alignment and the successful decoding of task-related labels at the final layer, we conduct a case study on SST-2 dataset to seek a semantic interpretation of the hidden states’ alignment dynamics.
@@
Surge in alignment concurs with the encoding of label-relevant semantics   First, we compute output alignment (logit lens accuracy) across all layers and compare its trajectory to directional alignment measures. As shown in Figure 4, their strong correlation indicates that output alignment rises in sync with directional alignment. This suggests that as alignment improves, label-relevant semantics are injected into hidden states via consecutive layer updates, enabling correct label decoding even at intermediate layers. To support this, we decode the top right singular vectors of hidden states near the transition point. Figure 4 (B) reveals a sharp semantic shift: while layer 20 produces irrelevant tokens, the SST-2 label token negative emerges as the second most probable token in layer 21—immediately post-transition. This confirms that post-transition layers elevate label-relevant directions in both the semantic and variance structure of the hidden states.
@@
Layer-wise pattern of semantic retention and filtering   To better understand how exactly layer updates after the phase transition encode relevant semantics into hidden states, we analyze the semantic information retained or removed by each layer. Given the SVDs of the centered hidden states at two consecutive layers $\bar{\bm{H}}_{l-1}=\bm{U}_{l-1}\bm{\Sigma}_{l-1}\bm{V}_{l-1}^{\top}$ and $\bar{\bm{H}}_{l}=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{\top}$, we compute the projection of the previous layer’s singular directions onto the next layer’s as $\bm{V}_{l-1}^{\top}\bm{V}_{l}\bm{V}_{l}^{\top}\in\mathbb{R}^{n\times d}$. The rows of this matrix with larger norms are the principal directions of $\bm{H}_{l-1}$ that are retained in $\bm{H}_{l}$, while those with smaller norms indicate filtered directions. By decoding these directions, we identify the semantic information retained or removed by each layer. The results are reported in Figure 4 (C). In most layers after the phase transition, label-related semantic directions are consistently retained and unrelated directions are filtered, leading to a progressive amplification of label-relevant information in the hidden states and a suppression of the irrelevant semantics. However, in the final layers, this pattern reverses: label-related directions are filtered, and label-irrelevant semantics are reintroduced. This shift corresponds to the slight drop in output alignment in the final layers in Figure 4 (A), and aligns with prior findings that late layers in LLMs tend to encode high-frequency, semantically uninformative tokens [Sharma et al., 2023], which may interfere with the expression of task-relevant content. The same phenomenon also occurs in other datasets as presented in Subsection F.4.
@@
Figure 5: Accuracy gains of rank-10 denoising. See Subsection F.5 for results in other settings.
@@
Low-rank denoising enhances label-relevant semantics   To better understand the retention-and-filtering process, we test whether filtering can be accelerated—and label-relevant semantics enhanced—by denoising the hidden states. Specifically, we apply a rank-10 SVD approximation to the centered hidden states $\bar{H}$, retaining only the top 10 directions of variation, and add the mean back. This denoising should theoretically remove less informative components and increase the prominence of label-relevant semantics. We then feed the denoised hidden states into the unembedding layer and measure output accuracy. As shown in Figure 5, low-rank denoising yields accuracy gains exceeding 10%, especially in early post-transition layers where irrelevant semantics are just beginning to be filtered. These results validate low-rank denoising and support semantic retention and filtering as a core mechanism of post-transition layers.
@@
Task Vector properties of post-transition hidden states   As visualized by Figure 1 (D), the simultaneous enhancement of both alignment properties and label-relevant semantics after the phase transition qualifies ICL query hidden states in middle-to-late layers as task vectors [Hendel et al., 2023]. Due to their improved alignment properties, steering or replacing zero-shot hidden states with ICL hidden states can align zero-shot hidden states better with the label unembedding vectors. Similarly, modifying zero-shot hidden states with ICL hidden states, which are label-informative, makes them more suited for the prediction of the correct labels.
@@ 5.3 Significance of critical attention heads
For each dataset and model, we rank all attention heads by their PTH and IH scores, which quantify how strongly they exhibit corresponding attention patterns on test prompts. We then ablate (zero out) the outputs of the top 10% identified IHs and PTHs and measure the resulting effects on hidden state separability and alignment across layers as well as the output accuracy. As a control, we repeat the procedure with an equal number of randomly selected heads (excluding the top PTH/IH heads). Details on identifying PTHs and IHs are provided in Appendix G.
@@
PTHs induce separability, IHs induce alignment   The results in Figure 6 confirm that PTHs and IHs are crucial not only for verbalizing ICL outputs [Cho et al., 2025a] but also for structuring hidden states both separability-wise and alignment-wise. Ablating random attention heads has a negligible effect (as shown by overlapping black and gray curves). In contrast, ablating PTHs or IHs significantly alters the measures, but with distinct effects. Ablating PTHs substantially reduces separability and also influences mean-based alignment (and thus composite alignment) which concerns the separation margin of the label clusters, but has a minimal impact on other alignment measures. Conversely, ablating IHs leaves separability largely intact but severely disrupts all alignment measures, and has a far greater impact on output alignment as measured by the accuracy.
Figure 7: Average effect across datasets of steering zero-shot hidden states using task vectors created from IH outputs.
@@
These results not only support the presence of a two-stage ICL process—initially driven by separability, then by alignment—but also clarify its mechanism. PTHs, concentrated in early layers [Cho et al., 2025a], enhance separability by attending to query tokens and encoding the distinct semantics of different queries into the query hidden states. IHs, which appear after a critical depth (30–40% of total layers [Halawi et al., 2024]), then copy demonstration label embeddings into the query hidden states through their outputs into the residual stream and improve their alignment with label unembedding directions. Moreover, in Subsection H.2 we show that PTH and IH contribute to ICL with the same mechanism even if the ground-truth label of query, which was considered a requirement for IH [Cho et al., 2025a], is not present in demonstrations.
@@
IH outputs as task vectors   Our findings also clarify why IH outputs serve as effective task vectors, as observed by Todd et al. [2024]. Since IH outputs can improve the alignment qualities of the hidden states, they can (1) boost zero-shot accuracy when added to zero-shot hidden states, and (2) decode label-related tokens, as they align with label unembedding directions. To validate this conclusion on our classification datasets (not considered in the original study), we extract task vectors from the outputs of the identified IHs and add them to layer-30 zero-shot hidden states of Llama2-70B (see Appendix I for details). The substantial accuracy gains reported in Figure 7 support this conclusion.
@@ 6 Conclusions and Limitations
Conclusions   This work unifies two major perspectives on in-context learning (ICL)—the roles of special attention heads and task vectors—within a geometric framework centered on the separability and alignment of hidden states. We theoretically show that these two properties fully determine ICL classification accuracy and empirically demonstrate that ICL improves accuracy over zero-shot primarily by enhancing alignment, explaining why ICL query hidden states serve effectively as task vectors. Our analysis reveals a phase transition: early layers boost separability, while later layers refine alignment with label unembedding vectors by filtering out label-irrelevant semantics and preserving task-relevant directions, which coincides with the emergence of label semantics in the hidden states. By zooming into layers, we identify the complementary roles of PTHs and IHs: PTHs, prevalent in early layers, enhance separability, while IHs, emerging post-transition, promote alignment—clarifying both the phase transition and the effectiveness of IH outputs as task vectors.
Limitations   Our framework is tailored to classification and may not extend to generation or multi-token outputs. In addition, our analysis focuses on inference-time behavior in pretrained models. Future work should investigate how separability and alignment emerge during training—e.g., at what stage do models begin to separate queries by label, and when does the phase transition arise?
@@ Appendix A Proofs related to Theorem 1
$\leq\sup_{\bm{u}\in R^{d},||\bm{u}||_{2}=1}\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}(\bm{u}^{T}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}(\bm{u}^{T}\bm{h}_{i}<0))$ supremum (7)
@@ Appendix B Generalization of Theorem 1
because for each ${\mathbb{N}}_{j}$, the term $\mathbbm{1}(y_{i}=\mathop{\arg\max}\limits_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ appears $(m-1)$ times in the summation for each $i\in{\mathbb{N}}_{j}$. Hence, combining everything, we have
@@ Appendix D Implementation details
Datasets We use the official huggingface implementation of all datasets except for the Famous People dataset crafted by ourselves. The Famous People dataset is generated using ChatGPT-4o [OpenAI et al., 2024]. The dataset has 180 datapoints, each is the name of a famous person with a profession label from one of six categories, which is one of the following six: actor, politician, singer, scientist, writer, athlete. The dataset is balanced as there are 30 data points for each label.
@@
For the label mapping experiment in Subsection 5.1. We use the mappings listed in Table 2 to map the ground-truth labels of each dataset to symbols.
@@
TREC abbreviation / entity / description / human / location / $\rightarrow$ @/#/\$/&/*
@@
Appendix E Supplementary materials for Subsection 5.1
E.1 Replication of Figure 2 for other models
@@
In Figure 8-13, we provide the visualizations of the experiments presented in Figure 2 for other models, from which similar conclusions to those in Subsection 5.1 can be drawn. Alignment measures explain the accuracy difference between ICL and zero-shot across models, and the layer-wise trends of the measures explain the phase transition pattern for all models.
E.2 Replication of Figure 3 for other models
@@
In Figure 14-19 we provide the visualizations of the experiments presented in Figure 3 for other models to establish the generality of our conclusions. It is clear that the phase transition pattern is evident in different ICL settings across models, and alignment measures capture accuracy differences across ICL settings for all models.
Appendix F Supplementary materials for Subsection 5.2
F.1 Replication of subplot (A) of Figure 4 for other datasets and models

### v2

@@ 5.2 Semantic interpretation of alignment dynamics
To understand how the surge in directional alignment after the phase transition leads to high output alignment and successful decoding of task-related labels at the final layer, we conduct a case study on the SST-2 dataset to interpret hidden-state alignment dynamics.
@@
Surge in alignment concurs with the encoding of label-relevant semantics  First, we compute output alignment (logit lens accuracy) across all layers and compare its trajectory to directional alignment measures. As shown in Figure˜5, their strong correlation indicates that output alignment rises in sync with directional alignment. This suggests that as alignment improves, label-relevant semantics are injected into hidden states via consecutive layer updates, enabling correct label decoding even at intermediate layers. To support this, we decode the top right singular vectors of hidden states near the transition point. Figure˜5 (B) reveals a sharp semantic shift: while layer 20 produces irrelevant tokens, the SST-2 label token negative emerges as the second most probable token in layer 21—immediately post-transition. This confirms that post-transition layers elevate label-relevant directions in both the semantic and variance structure of the hidden states.
Figure 5: Dynamics of alignment measures and semantics of post-transition hidden states. (A) Strong correlation between output and directional alignment; (B) Surge in alignment measures concur with encoding of label-related semantics; (C) Post-transition layers, except for the last ones, filter out unrelated semantics and retain the related ones. Refer to Appendix˜G for more semantics decoding cases.
@@
Layer-wise pattern of semantic retention and filtering  To better understand how exactly layer updates after the phase transition encode relevant semantics into hidden states, we analyze the semantic information retained or removed by each layer. Given the SVDs of the centered hidden states at two consecutive layers $\bar{\bm{H}}_{l-1}=\bm{U}_{l-1}\bm{\Sigma}_{l-1}\bm{V}_{l-1}^{\top}$ and $\bar{\bm{H}}_{l}=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{\top}$, we compute the projection of the previous layer’s singular directions onto the next layer’s as $\bm{V}_{l-1}^{\top}\bm{V}_{l}\bm{V}_{l}^{\top}\in\mathbb{R}^{n\times d}$. The rows of this matrix with larger norms are the principal directions of $\bm{H}_{l-1}$ that are retained in $\bm{H}_{l}$, while those with smaller norms indicate filtered directions. By decoding these directions, we identify the semantic information retained or removed by each layer. The results are reported in Figure˜5 (C). In most layers after the phase transition, label-related semantic directions are consistently retained, and unrelated directions are filtered, leading to a progressive amplification of label-relevant information in the hidden states and a suppression of the irrelevant semantics. However, in the final layers, this pattern reverses: label-related directions are filtered, and label-irrelevant semantics are reintroduced. This shift corresponds to the slight drop in output alignment in the final layers in Figure˜5 (A), and aligns with prior findings that late layers in LLMs tend to encode high-frequency, semantically uninformative tokens [Sharma et al., 2023], which may interfere with the expression of task-relevant content. The same phenomenon also occurs in other datasets as presented in Subsection˜G.4.
@@
Figure 6: Accuracy gains of rank-10 denoising. See Subsection˜G.5 for results in other settings.
@@
Low-rank denoising enhances label-relevant semantics  To better understand the retention-and-filtering process, we test whether filtering can be accelerated—and label-relevant semantics enhanced—by denoising the hidden states. Specifically, we apply a rank-10 SVD approximation to the centered hidden states $\bar{H}$, retaining only the top 10 directions of variation and adding the mean back. This denoising should theoretically remove less informative components and amplify label-relevant semantics. We then feed the denoised hidden states into the unembedding layer and measure output accuracy. As shown in Figure˜6, low-rank denoising yields accuracy gains exceeding 10%, especially in early post-transition layers where irrelevant semantics are just beginning to be filtered. These results validate low-rank denoising and support semantic retention and filtering as a core mechanism of post-transition layers. We explore the practical implication of this in Subsection˜E.2.
@@
Task vector properties of post-transition hidden states  As visualized by Figure˜1 (D), the simultaneous enhancement of both alignment properties and label-relevant semantics after the phase transition qualifies ICL query hidden states in middle-to-late layers as task vectors [Hendel et al., 2023]. Due to their improved alignment properties, steering or replacing zero-shot hidden states with ICL hidden states can align zero-shot hidden states better with the label unembedding vectors. Similarly, modifying zero-shot hidden states with ICL hidden states, which are label-informative, makes them more suited for the prediction of the correct labels.
@@ 5.3 Significance of critical attention heads
For each dataset and model, we rank all attention heads by their PTH and IH scores, which quantify how strongly they exhibit corresponding attention patterns on test prompts. We then ablate (zero out) the outputs of the top 10% identified IHs and PTHs and measure the resulting effects on hidden state separability and alignment across layers as well as the output accuracy. As a control, we repeat the procedure with an equal number of randomly selected heads (excluding the top PTH/IH heads). Details on identifying PTHs and IHs are provided in Appendix˜H.
@@
PTHs induce separability, IHs induce alignment  The results in Figure˜7 confirm that PTHs and IHs are crucial not only for verbalizing ICL outputs [Cho et al., 2025a] but also for structuring hidden states both separability-wise and alignment-wise. Ablating random attention heads has a negligible effect (as shown by overlapping black and gray curves). In contrast, ablating PTHs or IHs significantly alters the measures, but with distinct effects. Ablating PTHs substantially reduces separability and also influences mean-based alignment (and thus composite alignment) which concerns the separation margin of the label clusters, but has a minimal impact on other alignment measures. Conversely, ablating IHs leaves separability largely intact but severely disrupts all alignment measures, and has a far greater impact on output alignment as measured by the accuracy.
Table 1: Dataset-mean layer positions of top IH and PTH heads by percentage levels with statistically significant differences.
% IH mean PTH mean p-value
1% 40.203 33.379 $1.01\times 10^{-7}$
2% 40.884 36.459 $1.30\times 10^{-6}$
5% 42.238 36.422 $2.06\times 10^{-17}$
10% 41.535 35.850 $9.72\times 10^{-29}$
@@
These results not only confirm a two-stage ICL process—first driven by separability, then by alignment—but also clarify its mechanism. PTHs, concentrated in early layers [Cho et al., 2025a], enhance separability by attending to query tokens and encoding distinct semantics into hidden states. IHs, emerging after a critical depth (30–40% of total layers [Halawi et al., 2024]), copy demonstration label embeddings into query hidden states via the residual stream, improving alignment with label unembedding directions. Furthermore, in Subsection˜I.5, we show that PTHs and IHs contribute to ICL through the same mechanism even when the query’s ground-truth label—previously considered essential for IHs [Cho et al., 2025a]—is absent from demonstrations.
PTHs in early layers, IHs in late layers  To test whether PTHs and IHs align with our two-stage ICL characterization—inducing separability and alignment respectively—we examine their layer distributions. Specifically, we locate the top 1%, 2%, 5%, and 10% identified PTHs and IHs and perform Mann-Whitney U tests to assess differences across layers. As shown in Table 1, PTHs mainly occur in earlier layers, consistent with their role in inducing separability, while IHs appear later, with the differences being highly significant under the Mann-Whitney test.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 8: (A) PTH task vector increases separability after injection; (B) Injecting IH task vector recreates the phase-transition in alignment.
Geometric effects of IH and PTH outputs as task vectors  The ablation in Figure 7, though revealing how removing PTHs and IHs affects separability and alignment, does not show how they actively drive the geometric evolution of hidden states during inference. To address this, we follow Todd et al. [2024] and inject their outputs as task vectors into zero-shot hidden states at selected layers to observe induced metric changes. We construct task vectors using outputs of the top 10% IHs and PTHs as in Appendix˜J. We inject the PTH-derived vector at layer $l=2$ to track separability, and the IH-derived vector at $\frac{3}{8}L$ to track composite alignment. These layers correspond to where separability and alignment begin to shift markedly in Figure 2. Figure˜8(a) shows that injecting PTH outputs as task vectors boosts separability (the green line surpassing the magenta line before convergence; see Subsection˜I.2 for similar trends across models), while injecting IH task vectors restores the ICL-specific spike in alignment metrics in zero-shot hidden states.
Figure 9: Average effect across datasets of steering zero-shot hidden states using IH outputs as task vectors.
@@
IH outputs as task vectors  Since we have shown that IH outputs with ICL demonstrations can compensate for alignment deficits in hidden states, we have clarified how IH outputs as task vectors influence the forward computation of models: they steer hidden states to align with label-token unembeddings so these tokens can be decoded. This also explains why IH outputs themselves decode label-related tokens [Todd et al., 2024], as they must align with label unembedding directions to perform such steering. To validate this conclusion on our classification datasets (not considered in the original study), we extract task vectors from identified IH outputs and add them to layer-30 zero-shot hidden states of Llama2-70B. The substantial accuracy gains in Figure˜9 support this conclusion.
@@ 6 Conclusions and Limitations
Conclusions  This work unifies two major perspectives on in-context learning (ICL)—the roles of special attention heads and task vectors—within a geometric framework centered on separability and alignment of hidden states. We theoretically show that these two properties fully determine ICL classification accuracy and empirically demonstrate that ICL across classification and generation tasks improves accuracy over zero-shot mainly by enhancing alignment, explaining why ICL query hidden states serve effectively as task vectors. Our analysis reveals a phase transition: early layers boost separability, while later layers refine alignment with label unembedding vectors by filtering out label-irrelevant semantics and preserving task-relevant directions, coinciding with the emergence of label semantics in hidden states. Zooming into layers, we identify complementary roles of PTHs and IHs: PTHs in early layers enhance separability, while IHs emerging post-transition promote alignment—clarifying both the phase transition and the effectiveness of IH outputs as task vectors.
Limitations  Our analysis of the ICL hidden states of generation tasks is preliminary. In addition, our analysis focuses on inference-time behavior in pretrained models. Future work should investigate how the phase transition pattern of ICL hidden states emerges during training, which could shed light on the emergence of ICL capability.
Acknowledgments
This work was supported by JST FOREST Program (Grant Number JPMJFR232K, Japan) and the Nakajima Foundation. We used ABCI 3.0 provided by AIST and AIST Solutions with support from “ABCI 3.0 Development Acceleration Use”.
@@ Appendix A Proofs related to Theorem 1
[…]
@@ Appendix B Generalization of Theorem 1
because for each ${\mathbb{N}}_{j}$, the term $\mathbbm{1}(y_{i}=\mathop{\arg\max}\limits_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ appears (m1)(m-1) times in the summation for each $i\in{\mathbb{N}}_{j}$. Hence, combining everything, we have
@@ Appendix D Implementation details
Datasets  We use the official huggingface implementation of all datasets except for the Famous People dataset crafted by ourselves. The Famous People dataset is generated using ChatGPT-4o [OpenAI et al., 2024]. The dataset has 180 datapoints, each is the name of a famous person with a profession label from one of six categories: actor, politician, singer, scientist, writer, athlete. The dataset is balanced as there are 30 data points for each label.
@@
For the label mapping experiment in Subsection˜5.1. We use the mappings listed in Table 3 to map the ground-truth labels of each dataset to symbols.
@@
TREC abbreviation / entity / description / person / number / location / $\rightarrow$ @/#/!/\$/&/*
@@
Appendix E Practical implications of the mechanistic findings
We discuss in this section how our geometric characterization of the model’s inference process in the zero-shot or ICL setting can be converted to actionable and efficient methods to improve model performance.
E.1 Improving classification by capitalizing on hidden states separability
In Figure 2 we see the high separability of hidden states of prompts from classification tasks which surpasses the accuracy of the actual label prediction by far, irrespective of whether demonstrations are prepended to the labels. This implies that rather than direct decoding we can instead use the model as a feature extractor with which to train classifiers to perform the classification task, which is exactly how we evaluate the separability score. To demonstrate this, we train a logistic classifier on 50% of the last layer hidden states and evaluate the classification accuracy on the rest half. The results in Table 22 demonstrate that such a practice improves the classification accuracy with zero-shot prompts by over 70% across models, while only incurring the minimal cost of training an extra logistic classifier. Moreover, in Table 23 we see that training the classifier over 8-shot ICL hidden states also brings a solid $\sim$10% increase over the original ICL accuracy, which shows the generalizability of this lightweight method that leverages hidden states separability.
E.2 Unsupervised enhancement of classification performance through low-rank denoising
In Figure 6 we see that low-rank denoising of the ICL hidden states matrix reveals the task-related directions as its dominant directions and enhances its alignment properties. We can thus apply this technique to the final layer hidden states instead of the intermediate layer ones (which is the focus of Figure 6) to directly improve classification performance. In Table 24 we apply rank-5 denoising to final layer hidden states before feeding them into the unembedding layer. This method is different from the classifier-based method in that it assumes no knowledge of the ground-truth label of each hidden state. The dataset-average results demonstrate that the rank-5 denoising can indeed improve the classification performance of all models with the only exception of Gemma-7B.
Appendix F Supplementary materials for Subsection˜5.1
F.1 Replication of Figure 2 for other models
@@
In Figure 10-15, we provide the visualizations of the experiments presented in Figure 2 for other models, from which similar conclusions to those in Subsection˜5.1 can be drawn. Alignment measures explain the accuracy difference between ICL and zero-shot across models, and the layer-wise trends of the measures explain the phase transition pattern for all models.
F.2 Replication of Figure 3 for other models
@@
In Figure 16-21 we provide the visualizations of the experiments presented in Figure 3 for other models to establish the generality of our conclusions. It is clear that the phase transition pattern is evident in different ICL settings across models, and alignment measures capture accuracy differences across ICL settings for all models.
F.3 Replication of Figure 4 for other models
In Figure 53-58 we repeat the visualizations in Figure 4 for other models to examine where the generation task fits well into our framework. The results demonstrate that for other models, the phase-transition pattern in the composite alignment metric is specific to the 8-shot ICL hidden states, while ICL and zero-shot hidden states do not differ much in terms of the dynamics of the separability score.
Appendix G Supplementary materials for Subsection˜5.2
G.1 Replication of subplot (A) of Figure 5 for other datasets and models